# The efficiency of black hole formation via collisions in stellar systems:

### v1

@@ 3.1 Simulations data
Figure 1: Left: The black hole formation efficiency $\epsilon_{\mathrm{BH}}$ computed by Eq. 3 against the initial mass of the cluster $M_{\mathrm{i}}$ normalized by the critical mass $M_{\mathrm{crit}}$, calculated via Eq. 1. Different types of star clusters are represented by distinct colors: black symbols for $1^{\text{st}}$ star clusters (yellow symbols data from Reinoso et al. (2020), represent models with external potential, calculated via Eq. 2.), red symbols for GCs and blue symbols for NSCs. The final stellar mass is estimated neglecting mass loss via Eq. 4. Right: The black hole formation efficiency $\epsilon_{\mathrm{BH}}$ computed by Eq. 3 against the current mass of the cluster $M_{\mathrm{i}}$ normalized by the critical mass $M_{crit}$ calculated from Eq. 1, including an approximate correction for mass loss.
In simulations, the initial conditions are known, and the calculation of the critical mass $M_{\mathrm{crit}}$ and the ratio $M_{\mathrm{i}}/M_{\mathrm{crit}}$ is thus straightforward. Information about the final stellar mass in principle is part of the simulations but usually has not been reported in the literature. A first simple estimate can be obtained if the mass loss is neglected but considering the mass that has gone into the most massive object, implying
@@
The result of this calculation is given in the left panel of Fig. 1. While we found scatter in the plot, one can note clearly that higher efficiencies $\epsilon_{\mathrm{BH}}$ are obtained in all simulations with $M_{\mathrm{i}}/M_{\mathrm{crit}}\gtrsim 0.3$ with an increasing trend, the slope of this trend depending somewhat on the set of simulations being considered, with maximum efficiencies reaching around $40-50\%$. On the other hand for low values of $M_{\mathrm{i}}/M_{\mathrm{crit}}$, the relation appears essentially flat with efficiencies at best on the percent level.
In the left panel of Fig. 1, mass loss via escapers has been completely neglected, while in principle it is clear that it should be considered. We know however at least from some simulations how much mass has been lost during evolution. In appendix B, we employ the simulations of Vergara et al. (2021, 2023); Reinoso et al. (2020) to analyze how the ratio $M_{\mathrm{f}}/M_{\mathrm{i}}$ depends on $M_{\mathrm{i}}/M_{\mathrm{crit}}$. Such dependence can be expected for two reasons: a) if the mass of the most massive object increases with $M_{\mathrm{i}}/M_{\mathrm{crit}}$ as indicated via Fig. 1, this would imply such a relation even in the absence of escapers. b) escapers occur due to interactions within the stellar system, and a higher ratio of $M_{\mathrm{i}}/M_{\mathrm{crit}}$ implies that such interactions should be more frequent. In the appendix, we examine the dependence between these quantities both for the case when escapers are neglected (as when employing Eq. (4)) and when the full information from the simulations is being employed, and derive a correction factor $\alpha$ (Eq. 8) to estimate the expected mass loss due to escapers. We of course note that the exact amount of mass loss may depend on more factors including the adopted density profile and so the estimate arrived here can at best serve as an approximation.
Applying this correction factor may then lead to an improved estimate of the final stellar mass $M_{\mathrm{f}}$ and an improved calculation of $\epsilon_{\mathrm{BH}}$. The result of this calculation is given in the right panel of Fig. 1. The behaviour is qualitatively similar to the previous plot but overall appears more pronounced. The efficiency parameter $\epsilon_{\mathrm{BH}}$ now reaches up to $60\%$ for $M_{\mathrm{i}}/M_{\mathrm{crit}}\gtrsim 1$, though can also have more moderate values of the order $30-40\%$. An increase of the efficiency already appears visible from $M_{\mathrm{i}}/M_{\mathrm{crit}}\gtrsim 0.1$. For lower values of that ratio, the results appear very similar as before with efficiencies at best on the percent level. Therefore, the critical mass scale is probably not the only factor regulating the efficiency of collisions, although it still appears as a very important component regulating this process.
@@ 3.2 Observational data
For observations, as already noted above, we input the current conditions of the stellar system including $M_{\mathrm{f}}$ and $M_{\mathrm{BH}}$, allowing the direct calculation of $\epsilon_{\mathrm{BH}}$. However, it is to be noted that the initial properties for observational data are more uncertain as in the case of simulations, which are modeling ideal and isolated systems, real stellar clusters may not be isolated and some of them particularly the NSCs, may even experience recent or ongoing stellar formation and/or mergers with GCs. It is thus a relevant caveat to be taken that at least for the NSCs, the evolutionary history including mass evolution will be more complex as compared to the simulations. As discussed already in section 2.3, for NSCs we expect and assume an expansion during the evolution, while GCs are typically assumed to have approximately constant radii. However, we note that the half mass radius could increase by a factor of $2-3$ during the evolution (Gieles et al., 2010). The situation of UCDs is less clear and we will consider two possible scenarios as explained below. To estimate the initial mass for observations, we can consider as the sum of the current stellar mass and black hole mass within, under the assumption that at least its order of magnitude will be similar to $M_{\mathrm{i}}/M_{\mathrm{crit}}$ for simulations. The results are shown is given in Fig. 2. The scatter in this figure is somewhat increased, which due to the previous considerations perhaps is not surprising. Nonetheless, we note that for $M_{\mathrm{i}}/M_{\mathrm{crit}}\gtrsim 0.03$, $\epsilon_{\mathrm{BH}}$ is covering the entire range of efficiencies from the percent level up to $\sim 60\%$. On the other hand, for low values of $M_{\mathrm{i}}/M_{\mathrm{crit}}$, we only have efficiencies of up to $20\%$ at most. In the case of UCDs we explore one scenario (UCD1) where they are assumed to have fixed radius during their evolution and considering an external potential with $80\%$ of dark matter due to the observed mass to high ratios (e.g. Baumgardt & Mieske, 2008). We also explore another case (UCD2), more compatible with Fellhauer & Kroupa (2006), in the absence of dark matter. There, we assume that an initial expansion of the systems takes place, like in the case of NSCs. We find that particularly the case of UCD2 fits very well into the picture and displays a quite similar behaviour as the NSCs. The assumptions made in the case of UCD1 would somewhat enhance the scatter, while however not fundamentally changing the picture. We also note that high efficiencies close to $100\%$ are only found when the ratio of stellar mass over critical mass is above $1$, while otherwise typically it ranges from low efficiencies up to the order of $20\%$. We further note that we do not know the physics of how the massive black holes in these systems have formed, as it is at least conceivable that also gas physics should have been involved. The latter may contribute to adding scatter in the plot and increase the efficiencies above the expectation from a purely collisional channel. Particularly when considering the group of GCs, we note that they all behave very similarly, and in terms of the physics it is perhaps also the case that is most clear, as they should evolve essentially in isolation after their initial formation, while on the other hand, the NSCs are very likely the most complex systems we consider here, with possibly different formation histories, their contribution very considerably increases the scatter we find in this plot, and the UCDs might be somewhere in between GCs and NSCs judging from their respective scatter within the plot.
@@
To assess the sensitivity of this plot to the aforementioned assumptions, we have estimated $M_{\mathrm{i}}$ as the sum of the current stellar mass and mass of the black hole within. The critical mass range varies depending on the stellar system age and their initial properties. For NSCs, it typically falls within $10^{5}-10^{9}\leavevmode\nobreak\ \rm{\mathrm{M_{\odot}}}$, for GCs, it is between $10^{6}-10^{8}\leavevmode\nobreak\ \rm{\mathrm{M_{\odot}}}$, and for UCDs, it ranges from $10^{9}-10^{11}\leavevmode\nobreak\ \rm{\mathrm{M_{\odot}}}$ (UCD1) or $10^{7}-10^{9}\leavevmode\nobreak\ \rm{\mathrm{M_{\odot}}}$ (UCD2). The results given in Fig. 2 are in particular overall very similar to the left panel of Fig. 1. We note that for low values of $M_{\mathrm{i}}/M_{\mathrm{crit}}$, efficiencies can reach at best up to $20\%$, while for $M_{\mathrm{i}}/M_{\mathrm{crit}}\gtrsim 0.03$ the whole range of efficiencies up to $100\%$ is possible. Our qualitative conclusions are thus similar though the evolutionary history is more complex as compared to the simulations.
@@
Figure 2: The black hole formation efficiency $\epsilon_{BH}$ computed via Eq. 3 against the initial mass of the cluster normalized by the critical mass $M_{crit}$ (estimated from the current properties of the cluster), for observational data.
@@
We note three special cases. VCC1254 which introduces a higher scatter at $M_{\mathrm{i}}/M_{crit}\sim 10^{-2}$ with $\epsilon_{\mathrm{BH}}$ values around $50\%$. However, the black hole mass of VCC1254 is an upper limit. Besides having a larger nuclear mass-to-light ratio, which means that it must contain an older stellar population (Geha et al., 2002), contrarily to observations by Durrell (1997) and Stiavelli et al. (2001) that show the presence of young stars. Thus more precise measurements will lead to a more accurate determination of the mass and thus lower $\epsilon_{\mathrm{BH}}$. The UCDs M60-UCD1 and UCD320 show $\epsilon_{\mathrm{BH}}\approx 15-30\%$ at $M_{\mathrm{i}}/M_{crit}\sim 10^{-3}-10^{-2}$, respectively under the assumptions of UCD1. Compared to the stellar mass, its high black hole mass suggests that its origin could be a tidal core of a dwarf galaxy (Bassino et al., 1994; Bekki & Freeman, 2003; Pfeffer & Baumgardt, 2013; Seth et al., 2014; Afanasiev et al., 2018), which aligns with the case of UCD2. Finally, the scatter presented in some observations must be related to their complex formation history which is beyond the scope of this investigation.
@@
Finally we also note the case of $\mathrm{\omega}$ Cen (NGC 5139). This GC has been widely studied due to its unique characteristic of being the most massive GC of the Milky Way. Several studies have suggested that the stellar dynamics in $\mathrm{\omega}$ Cen can be explained due to the presence of an IMBH with a mass $\sim 4\times 10^{4}\leavevmode\nobreak\ \rm{\mathrm{M_{\odot}}}$ at the center (Noyola et al., 2008; Noyola & Baumgardt, 2011; Jalali et al., 2012; Baumgardt, 2017). Others argue that the stellar kinematics can be explained by the presence of a dense cluster of stellar mass black holes (Baumgardt et al., 2019; Zocchi et al., 2019). Other studies suggest that the presence of an IMBH is not necessary to explain the stellar dynamics (van der Marel & Anderson, 2010). In this investigation, we consider the first option to determine an upper limit of the efficiency. We note that for the other two options the black hole formation efficiency will be $\epsilon_{\mathrm{BH}}\simeq 0$, this result still aligns with the trend.
@@ 3.3 Comparative analysis of observational data and simulations data
So far we have independently analyzed the data from simulations and observations to check whether they support the concept of a critical mass that significantly determines the efficiency of collisions in dense stellar systems. We already noted the main differences regarding the typical data that are reported for the respective simulations and observations and developed the tools to at least approximately estimate the quantities that cannot be directly inferred. During the analysis of the observational data, we further encountered some of the systematic differences between simulations that usually model idealized and isolated stellar systems, while real stellar systems, in particular NSCs, are likely more complex due to their star formation histories as well as occasional mergers with GCs. On the other hand, the GCs appear to at least more closely resemble the behaviour of an isolated system as their star formation terminated very long ago, while UCDs are somewhat more uncertain as their origin is not known.
@@
Having these differences in mind, we aim to approximately assess what simulations and observations may have in common and which differences can be identified from the comparison. This may, on one hand, allow us to establish processes that play an important role in both observations and simulations, but also help to identify processes that are not yet included in simulations, which may play a possibly important role during their dynamical evolution.
@@
For this analysis here, in the context of the simulations, the mass loss will be estimated using a correction factor $\alpha$ (see Eq. 8) when calculating $\epsilon_{\mathrm{BH}}$. In the observations, we have all the information on the current stage. The result of this calculation is given in Fig. 3, where the top panel shows the data comparing all simulations and observations, while in the mid and bottom panels, we focus on data (simulations and observations) for NSCs and GCs, respectively. Note that in the mid panel, we add observations of UCDs, under the assumption of no dark matter but expansion (UCD2), while in the bottom panel, we add observations of UCDs, under the assumption of the presence of dark matter and no expansion (UCD1).
@@
In general, there is a broad consistency between the data from simulations and observations. Where the efficiencies are on the percent level when we consider $M_{\mathrm{i}}/M_{\mathrm{crit}}$ to be low, and only for $M_{\mathrm{i}}/M_{\mathrm{crit}}\gtrsim 0.1$ we find higher efficiencies up to of the order $100\%$, but with significant scatter. Part of the scatter arises from the NSCs for the reasons we were discussing above, i.e. the complex star formation histories and mergers with other stellar systems. If we consider globular cluster systems, simulations and observations appear very consistent and form a tight relation with efficiencies on the percent level at best. The UCDs are most uncertain while the case of UCD2 (no dark matter, expansion) fits best within the general trend, also the assumptions of UCD1 (dark matter, no expansion) fit into the general picture. We also emphasize that due to the unclear origin and the complex physics in these systems, a complete agreement could also be coincidental.
@@
Figure 3: The black hole formation efficiency $\epsilon_{\mathrm{BH}}$ computed by Eq. 3 against the current mass of the cluster $M_{\mathrm{i}}$ normalized by the critical mass $M_{crit}$ (Eq. 1). Top: for all observations and simulations. Middle: for observations and simulations of NSC, also include observations of UCD2. Bottom: for observations and simulations of GC, also include observations of UCD1.

### v2

@@ 3.1 Simulations data
Figure 1: Black hole formation efficiency, $\epsilon_{\mathrm{BH}}$, (left) computed by Eq. 3 against the initial mass of the cluster, $M_{\mathrm{i}}$, normalized by the critical mass, $M_{\mathrm{crit}}$, calculated via Eq. 1. Different types of star clusters are represented by distinct colors: black symbols for the first star clusters (yellow symbols data from Reinoso et al. (2020), represent models with external potential, calculated via Eq. 2.), red symbols for GCs and blue symbols for NSCs. The final stellar mass is estimated neglecting mass loss via Eq. 4. Right: The black hole formation efficiency $\epsilon_{\mathrm{BH}}$ computed by Eq. 3 against the current mass of the cluster, $M_{\mathrm{i}}$, normalized by the critical mass, $M_{crit}$, calculated from Eq. 1, including an approximate correction for mass loss.
In simulations, the initial conditions are known, and the calculation of the critical mass, $M_{\mathrm{crit}}$, and the ratio, $M_{\mathrm{i}}/M_{\mathrm{crit}}$, is thus straightforward. Information about the final stellar mass is, in principle, part of the simulations, but has not usually been reported in the literature. A first simple estimate can be obtained if the mass loss is neglected, but still considering the mass that has gone into the most massive object, implying:
@@
The result of this calculation is given in the left panel of Fig. 1. While we found scatter in the plot, we can clearly note that higher efficiencies, $\epsilon_{\mathrm{BH}}$, are obtained in all simulations with $M_{\mathrm{i}}/M_{\mathrm{crit}}\gtrsim 0.3$ with an increasing trend, the slope of this trend depending somewhat on the set of simulations being considered, with maximum efficiencies reaching around $40-50\%$. On the other hand, for low values of $M_{\mathrm{i}}/M_{\mathrm{crit}}$, the relation appears essentially flat with efficiencies (at best) on the level of a few percent.
In the left panel of Fig. 1, mass loss via escapers has been completely neglected, while (in principle) it is clear that it should be considered. We know however at least from some simulations how much mass has been lost during evolution. In Appendix B, we employ the simulations of Vergara et al. (2021, 2023); Reinoso et al. (2020) to analyze how the ratio $M_{\mathrm{f}}/M_{\mathrm{i}}$ depends on $M_{\mathrm{i}}/M_{\mathrm{crit}}$. Such a dependence can be expected for two reasons: a) if the mass of the most massive object increases with $M_{\mathrm{i}}/M_{\mathrm{crit}}$ as indicated via Fig. 1, this would imply such a relation even in the absence of escapers; b) escapers occur due to interactions within the stellar system and a higher ratio of $M_{\mathrm{i}}/M_{\mathrm{crit}}$ implies that such interactions should be more frequent. In the appendix, we examine the dependence between these quantities both for the case when escapers are neglected (as when employing Eq. (4)) and when the full information from the simulations is being employed. We derived a correction factor $\alpha$ (Eq. 8) to estimate the expected mass loss due to escapers. Of course, we note that the exact volume of mass loss may depend on more factors including the adopted density profile and so the estimate arrived here can at best serve as an approximation.
Applying this correction factor may then lead to an improved estimate of the final stellar mass, $M_{\mathrm{f}}$, and an improved calculation of $\epsilon_{\mathrm{BH}}$. The result of this calculation is given in the right panel of Fig. 1. The behavior is qualitatively similar to what is seen in the previous plot, but overall it appears more pronounced. The efficiency parameter $\epsilon_{\mathrm{BH}}$ now reaches up to $60\%$ for $M_{\mathrm{i}}/M_{\mathrm{crit}}\gtrsim 1$, although this may also have more moderate values, on the order of $30-40\%$. An increase in the efficiency already appears visible from $M_{\mathrm{i}}/M_{\mathrm{crit}}\gtrsim 0.1$. For lower values of that ratio, the results appear very similar as before with efficiencies at best on the percent level. Therefore, the critical mass scale is probably not the only factor regulating the efficiency of collisions, although it still appears to be a very important component regulating this process.
@@ 3.2 Observational data
For observations, as already noted above, we input the current conditions of the stellar system including $M_{\mathrm{f}}$ and $M_{\mathrm{BH}}$, allowing for the direct calculation of $\epsilon_{\mathrm{BH}}$. However, it is to be noted that the initial properties for observational data are more uncertain as in the case of simulations, which are meant to model ideal and isolated systems; real stellar clusters may not be isolated and some of them (particularly the NSCs) may even experience recent or ongoing stellar formation and/or mergers with GCs. It is thus a relevant caveat that needs to be taken (at least for the NSCs), namely, that the evolutionary history, including mass evolution, will be more complex as compared to the simulations. As discussed already in Section 2.3, for NSCs, we expect and assume an expansion during the evolution, while GCs are typically assumed to have approximately constant radii. However, we note that the half-mass radius could increase by a factor of $2-3$ during the evolution (Gieles et al., 2010). The situation of UCDs is less clear and thus, we consider two possible scenarios below. To estimate the initial mass for observations, we can consider it as the sum of the current stellar mass and black hole mass it contains, under the assumption that at least its order of magnitude will be similar to $M_{\mathrm{i}}/M_{\mathrm{crit}}$ for simulations. The results are shown is given in Fig. 2. The scatter in this figure is somewhat increased, which given the previous considerations is perhaps unsurprising. Nonetheless, we note that for $M_{\mathrm{i}}/M_{\mathrm{crit}}\gtrsim 0.03$, $\epsilon_{\mathrm{BH}}$ is covering the entire range of efficiencies from the percent level up to $\sim 60\%$. On the other hand, for low values of $M_{\mathrm{i}}/M_{\mathrm{crit}}$, we only have efficiencies of up to $20\%$ at most. In the case of UCDs, we explore one scenario (UCD1), where they are assumed to have fixed radius during their evolution and considering an external potential with $80\%$ of dark matter due to the observed mass to high ratios (e.g., Baumgardt & Mieske, 2008). We also explore another case (UCD2), more compatible with Fellhauer & Kroupa (2006), in the absence of dark matter. There, we assume that an initial expansion of the systems takes place, as in the case of NSCs. We find that the case of UCD2 in particular fits very well into the picture and displays a quite similar behavior as the NSCs. The assumptions made in the case of UCD1 would somewhat enhance the scatter, whereas it does not fundamentally change the picture. We also note that high efficiencies close to $100\%$ are only found when the ratio of stellar mass over critical mass is above $1$; otherwise, it typically ranges from low efficiencies up to the order of $20\%$. We further note that we do not know the physics of how the massive black holes in these systems have formed, as it is at least conceivable that also gas physics should have been involved. The latter may contribute to adding scatter in the plot and increase the efficiencies above the expectation from a purely collisional channel. In particular, when considering the group of GCs, we note that they all behave very similarly; in terms of the physics, it is perhaps also the case that is most clear, as they are expected to evolve essentially in isolation after their initial formation. On the other hand, the NSCs are very likely the most complex systems we consider here, with possibly different formation histories, their contribution very considerably increases the scatter we find in this plot. Thus, the UCDs might be somewhere in between GCs and NSCs judging from their respective scatter within the plot.
@@
To assess the sensitivity of this plot to the aforementioned assumptions, we estimated $M_{\mathrm{i}}$ as the sum of the current stellar mass and mass of the black hole within. The critical mass range varies depending on the stellar system age and their initial properties. For NSCs, it typically falls within $10^{5}-10^{9}\leavevmode\nobreak\ \rm{\mathrm{M_{\odot}}}$, for GCs, it is between $10^{6}-10^{8}\leavevmode\nobreak\ \rm{\mathrm{M_{\odot}}}$, and for UCDs, it ranges from $10^{9}-10^{11}\leavevmode\nobreak\ \rm{\mathrm{M_{\odot}}}$ (UCD1) or $10^{7}-10^{9}\leavevmode\nobreak\ \rm{\mathrm{M_{\odot}}}$ (UCD2). The results given in Fig. 2 are in particular overall very similar to the left panel of Fig. 1. We note that for low values of $M_{\mathrm{i}}/M_{\mathrm{crit}}$, efficiencies can reach at best up to $20\%$, while for $M_{\mathrm{i}}/M_{\mathrm{crit}}\gtrsim 0.03$ the whole range of efficiencies up to $100\%$ is possible. Our qualitative conclusions are thus similar though the evolutionary history is more complex as compared to the simulations.
@@
Figure 2: Black hole formation efficiency, $\epsilon_{BH}$, computed via Eq. 3 against the initial mass of the cluster normalized by the critical mass, $M_{crit}$ (estimated from the current properties of the cluster), for observational data.
@@
We note three special cases: VCC1254, which introduces a higher scatter at $M_{\mathrm{i}}/M_{crit}\sim 10^{-2}$, with $\epsilon_{\mathrm{BH}}$ values around $50\%$. However, the black hole mass of VCC1254 is an upper limit. Besides having a larger nuclear mass-to-light ratio, which means that it must contain an older stellar population (Geha et al., 2002), in contrast to observations by Durrell (1997) and Stiavelli et al. (2001), which show the presence of young stars. Thus more precise measurements will lead to a more accurate determination of the mass and thus lower $\epsilon_{\mathrm{BH}}$. The UCDs M60-UCD1 and UCD320 show $\epsilon_{\mathrm{BH}}\approx 15-30\%$ at $M_{\mathrm{i}}/M_{crit}\sim 10^{-3}-10^{-2}$, respectively, under the assumptions of UCD1. Compared to the stellar mass, its high black hole mass suggests that its origin could be a tidal core of a dwarf galaxy (Bassino et al., 1994; Bekki & Freeman, 2003; Pfeffer & Baumgardt, 2013; Seth et al., 2014; Afanasiev et al., 2018), which aligns with the case of UCD2. Finally, the scatter presented in some observations must be related to their complex formation history, which is beyond the scope of this investigation.
@@
Finally we also note the case of $\mathrm{\omega}$ Cen (NGC 5139). This GC has been widely studied due to its unique characteristic of being the most massive GC of the Milky Way. Several studies have suggested that the stellar dynamics in $\mathrm{\omega}$ Cen could be explained due to the presence of an IMBH with a mass $\sim 4\times 10^{4}\leavevmode\nobreak\ \rm{\mathrm{M_{\odot}}}$ at the center (Noyola et al., 2008; Noyola & Baumgardt, 2011; Jalali et al., 2012; Baumgardt, 2017). Others have argued that the stellar kinematics can be explained by the presence of a dense cluster of stellar-mass black holes (Baumgardt et al., 2019; Zocchi et al., 2019). Other studies suggest that the presence of an IMBH is not necessary to explain the stellar dynamics (van der Marel & Anderson, 2010). In this investigation, we consider the first option to determine an upper limit for the efficiency. We note that for the other two options, the black hole formation efficiency will be $\epsilon_{\mathrm{BH}}\simeq 0$, as this result still aligns with the trend.
@@ 3.3 Comparative analysis of observational data and simulations data
So far, we have independently analyzed the data from simulations and observations to check whether they support the concept of a critical mass that significantly determines the efficiency of collisions in dense stellar systems. We have noted the main differences regarding the typical data that are reported for the respective simulations and observations and developed the tools to at least approximately estimate the quantities that cannot be directly inferred. During the analysis of the observational data, we further encountered some of the systematic differences between simulations that usually model idealized and isolated stellar systems. However, real stellar systems, in particular NSCs, are likely more complex due to their star formation histories as well as occasional mergers with GCs. On the other hand, the GCs appear to at least more closely resemble the behavior of an isolated system as their star formation terminated very long ago, while UCDs are somewhat more uncertain as their origin is not known.
@@
With these differences in mind, we aim to approximately assess what simulations and observations may have in common and which differences can be identified from the comparison. This may, on one hand, allow us to establish processes that play an important role in both observations and simulations, but also help to identify processes that are not yet included in simulations, which may play a possibly important role during their dynamical evolution.
@@
For this analysis here, in the context of the simulations, the mass loss will be estimated using a correction factor, $\alpha,$ (see Eq. 8) when calculating $\epsilon_{\mathrm{BH}}$. In the observations, we have all the information on the current stage. The result of this calculation is given in Fig. 3, where the top panel shows the data comparing all simulations and observations, while in the mid and bottom panels, we focus on data (simulations and observations) for NSCs and GCs, respectively. We note that in the mid panel, we add observations of UCDs, under the assumption of no dark matter but expansion (UCD2), while in the bottom panel, we add observations of UCDs, under the assumption of the presence of dark matter and no expansion (UCD1).
@@
In general, there is a broad consistency between the data from simulations and observations. Where the efficiencies are on the percent level when we consider $M_{\mathrm{i}}/M_{\mathrm{crit}}$ to be low, and only for $M_{\mathrm{i}}/M_{\mathrm{crit}}\gtrsim 0.1$ we find higher efficiencies up to of the order $100\%$, but with significant scatter. Part of the scatter arises from the NSCs for the reasons we were discussing above, namely, the complex star formation histories and mergers with other stellar systems. If we consider globular cluster systems, the simulations and observations appear very consistent and form a tight relation with efficiencies on the percent level at best. The UCDs are most uncertain, while the case of UCD2 (no dark matter, including expansion) fits best within the general trend, also the assumptions of UCD1 (dark matter, no expansion) fit into the general picture. We also emphasize that due to the unclear origin and the complex physics in these systems, a complete agreement could also be coincidental.
@@
Figure 3: Black hole formation efficiency $\epsilon_{\mathrm{BH}}$ computed by Eq. 3 against the current mass of the cluster $M_{\mathrm{i}}$ normalized by the critical mass $M_{crit}$ (Eq. 1) for all observations and simulations (top); for observations and simulations of NSC, also including observations of UCD2 (middle); and for observations and simulations of GC, also including observations of UCD1 (bottom).